A small coin, initially at rest, begins falling. If the clock starts when the coin begins to fall, what is the magnitude of the coin's displacement between ?1=0.219 s and ?2=0.984 s?

Δ?=___________?

15) You are an engineer in charge of designing a new generation of elevators for a prospective upgrade to the Empire State Building. Before the state legislature votes on funding for the project, they would like you to prepare a report on the benefits of upgrading the elevators. One of the numbers that they have requested is the time it will take the elevator to go from the ground floor to the 102nd floor observatory. They are unlikely to approve the project unless the new elevators make the trip much faster than the old elevators.

If state law mandates that elevators cannot accelerate more than 1.50 m/s2 or travel faster than 12.3 m/s, what is the minimum time in which an elevator can travel the 373 m from the ground floor to the observatory floor?

minimum travel time:__________
Answer #1

14.

Solution-

Here we see that the acting force is only gravitational force

So using the Newtons law we have

=>ma = mg

Here m =The mass of the object

g =The acceleration due to gravity
=>a = g
So a= 9.8m/s^2

The initial velocity u = 0, and let us assume the displacement be s1 and s2

Using the formula

s = ut + 0.5at^2
u = 0 , a = g = 9.8 m/s^2
s1 = 0.5*9.8*(0.219)^2
s1 = 0.235 m

and
at t = 0.984
s2 = 0.5*9.8*(0.984)^2
s2 = 4.744 m


Therefore between both times the displacement is
=>s2-s1 =4.744-0.235

= 4.509 m